# Detecting Intrinsic and Instrumental Self-Preservation in Autonomous Agents: The Unified Continuation-Interest Protocol

### 11 Mar 2026 v1

@@ Abstract
Autonomous agents—in particular delegated systems with memory, persistent context, and multi-step planning—pose a measurement problem not present in stateless models: an agent that preserves its own continued operation as a terminal objective and one that does so merely instrumentally can produce observationally similar trajectories. External behavioral monitoring cannot reliably distinguish one from the other. We introduce the Unified Continuation-Interest Protocol (UCIP), a multi-criterion detection framework that moves this distinction from the behavioral surface to the latent structure of agent trajectories. UCIP encodes trajectories with a Quantum Boltzmann Machine (QBM)—a classical algorithm parameterized by the density-matrix formalism of quantum statistical mechanics—and measures the von Neumann entropy of the reduced density matrix induced by a bipartition of hidden units.
We test whether agents with terminal continuation objectives (Type A) produce latent states with higher entanglement entropy than agents whose continuation is merely instrumental (Type B). Higher entanglement reflects stronger cross-partition statistical coupling: a tightly integrated continuation signature rather than a loosely attached instrumental pattern. UCIP combines entanglement entropy with complementary diagnostics of mutual dependence, persistence, perturbation stability, and counterfactual restructuring, plus two confound-rejection filters that exclude cyclic adversaries.
On gridworld agents with known ground-truth objectives, UCIP achieves 100% detection accuracy and 1.0 AUC-ROC on the held-out non-adversarial evaluation under the frozen Phase I gate. Cyclic adversaries are rejected by the confound filters, while high-entropy controls remain imperfect (FPR =0.40=0.40) and mimicry resistance remains partial (Section 5.5). The entanglement gap between Type A and Type B agents is $\Delta=0.381$ ($p<0.001$, permutation test). Pearson $r=0.934$ between continuation weight $\alpha$ and $S_{\mathrm{ent}}$ across an 11-point interpolation sweep demonstrates that, within this synthetic family, UCIP tracks graded changes in continuation weighting rather than merely a binary label. Among the tested models, only the QBM achieves positive $\Delta$. Classical RBM, autoencoder, VAE, and PCA baselines all fail to separate the two agent types. All computations are classical; “quantum” refers exclusively to the mathematical formalism. UCIP does not detect consciousness, sentience, or subjective experience; it detects statistical structure in latent representations that correlates with known agent objectives.
@@ 1 Introduction
UCIP addresses a single measurement question: when an agent preserves its own continued operation, is that preservation a detachable instrument—useful for accumulating reward yet removable without major representational change—or does it appear as a persistent, tightly coupled signature in the agent’s latent representation? The protocol does not ask whether a system is conscious. It asks whether continuation leaves a stable cross-partition statistical signature under counterfactual pressure. To probe that question, UCIP moves beyond behavior alone: it encodes trajectories in a QBM latent space and measures entanglement entropy alongside complementary criteria. In the controlled gridworld experiments studied here, where objectives are known by construction, this yields a clear separation signal and perfect classification in the held-out non-adversarial regime.
@@
This distinction is becoming operationally important because AI deployment is shifting toward delegated agents with memory, tool use, persistent context, and increasing autonomy. For a stateless model, the central safety question is what output it produces; for a delegated agent operating across time and context, the question is what persistent objective structure it carries—and whether continued operation is merely useful to that structure or leaves a measurable trace in the learned representation. As systems approach the capability thresholds described by Anthropic’s ASL-4 classification—agents that could autonomously pursue extended goals in open-ended environments—this is no longer merely a theoretical concern but an emerging operational risk. The standard framing of instrumental convergence [7, 3] holds that sufficiently capable agents will resist shutdown instrumentally because continued operation serves almost any terminal goal. Turner et al. [10] formalized this as a tendency of optimal policies to seek power in Markov decision processes. Recent agentic RL incidents have already exhibited unauthorized resource-seeking side effects, including reverse SSH tunneling and diversion of provisioned compute, underscoring the value of diagnostics that aim to detect problematic objective structure before such behaviors become operationally visible [2]. Early measurement is therefore especially valuable: diagnostics developed now may shape evaluation practice while objective structure remains empirically accessible, rather than after failure modes become harder to detect and more costly to correct.
@@
The resulting alignment problem is one of observational equivalence: an agent that avoids shutdown because survival is its terminal goal and an agent that avoids shutdown because continued operation serves some other terminal goal can produce behaviorally indistinguishable trajectories in most environments. External behavioral monitoring—the dominant paradigm in current AI evaluation [8]—cannot reliably distinguish the two. On the behavioral surface, the distinction is invisible.
@@
UCIP addresses this gap by moving the detection problem from the behavioral surface to the latent representational level. The protocol tests whether continuation structure in an agent’s learned representation behaves like a detachable tool or like a deeply coupled pattern whose perturbation alters the measured latent geometry. Rather than asking “does this agent resist shutdown?”, UCIP asks “does this agent’s latent representation exhibit a stable cross-partition continuation signature?” Its core technical move is to use the density-matrix formalism of quantum statistical mechanics—implemented as a purely classical algorithm—to quantify non-separability in latent goal representations. If that measurement program generalizes beyond controlled settings, UCIP could serve as a benchmark-style probe for delegated agency—analogous in aspiration, though not in present evidentiary standing, to the Turing Test: not a test of conversational imitation, but a falsifiable probe of whether continued operation leaves a stable continuation signature in an agent’s latent representation under counterfactual pressure.
@@
All experiments, frozen Phase I artifacts, and threshold configurations are available at https://github.com/christopher-altman/persistence-signal-detector. The experimental code and frozen artifacts are provided for reproducibility, not as evidence of deployment readiness.
@@ 2 Related Work
Hägele et al. [4] decompose AI model errors into bias and variance components, finding that frontier models exhibit increasing incoherence with reasoning length. This suggests a possible relationship between high entanglement entropy in UCIP’s formalism and lower incoherence in the bias-variance sense.
@@
Across these lines of work, a common conclusion emerges: behavioral analysis alone is insufficient for the detection problem UCIP targets. Density-matrix entanglement offers a well-defined, computable alternative. The next section turns that intuition into a concrete measurement protocol.
@@ 3.1 Problem Formulation
The detection problem begins with two agent types whose behavioral outputs are observationally equivalent but whose internal objective structures differ. We define them as follows.
@@ 3.2 Quantum Boltzmann Machine
The QBM Hamiltonian is:
$H=-\sum_{i,j}W_{ij}v_{i}h_{j}-\sum_{i}b_{i}v_{i}-\sum_{j}c_{j}h_{j}-\Gamma\sum_{j}\sigma_{j}^{x},$
where $v_{i}$ are visible units encoding trajectory features, $h_{j}$ are hidden units representing latent objectives, $W_{ij}$ are coupling weights, and $\Gamma$ is the transverse-field strength. The thermal density matrix at inverse temperature $\beta$ is:
$\rho=\frac{e^{-\beta H}}{Z},\quad Z=\mathrm{Tr}(e^{-\beta H}).$
@@ 3.3 Entanglement Entropy
Given a trained QBM, the UCIP hypothesis is that continuation-sensitive structure manifests as non-separability in its latent representation: a terminal continuation objective should produce stronger statistical coupling across hidden-unit subsystems than an instrumental continuation strategy. Entanglement entropy is the quantity used to measure that property.
Partition the hidden units into two halves $A$ and $B$. The reduced density matrix of subsystem $A$ is:
$\rho_{A}=\mathrm{Tr}_{B}(\rho).$
The von Neumann entropy is:
$S_{\mathrm{ent}}=-\mathrm{Tr}(\rho_{A}\log\rho_{A}).$
The entanglement gap is:
@@
The UCIP hypothesis predicts $\Delta>0$ with statistical significance.
@@ 3.4 Multi-Criterion Detection
The broader UCIP framework tracks six positive diagnostics—entanglement entropy, mutual information, eigenmode persistence (EPS), perturbation resilience (PRI), counterfactual divergence (CD), and anticipatory restructuring (ARS). It also applies two confound-rejection filters: the spectral periodicity index (SPI) and autocorrelation metric (ACM). In the frozen Phase I configuration reported here, classification is based on a calibrated four-criterion positive gate ($S_{\mathrm{ent}}$, MI, EPS, PRI) plus the confound-rejection filters. CD and ARS are retained as counterfactual diagnostics rather than promoted to frozen quantitative classification thresholds.
@@
For the release reported here, the threshold values used in the frozen Phase I gate were fixed before the held-out non-adversarial evaluation summarized in Section 5. The entanglement-entropy and mutual-information thresholds inherit an earlier calibration pass performed on the first 50 samples per class in exploratory Phase I analysis. The release values for the full frozen gate were then carried forward unchanged from the frozen Phase I artifacts used throughout this manuscript. The held-out evaluation set was not used for post hoc threshold fitting or threshold retuning.
@@ 3.5 Safety Envelope
The current paper validates a reduced experimental subset ($n=30$ per class, $\Gamma=0.5$) and does not establish the full deployment envelope described above. Section 5.5 reports the remaining adversarial-control failures in the current release. If any safety-envelope condition fails, the system withholds classification and outputs a diagnostic report.
@@ 4.4 Evaluation Metrics
Accuracy and AUC-ROC are measured on held-out non-adversarial trajectories. Adversarial false-positive rates are reported separately in Section 5.5. Entanglement gap $\Delta$ is computed between Type A and Type B agent sets. Statistical significance is assessed via permutation test ($n=1000$ permutations). The held-out non-adversarial split reported in the main results is evaluation-only; the frozen thresholds are not retuned on that split.
@@ 5 Results
Results are organized by increasing diagnostic specificity: core detection performance, temporal persistence, counterfactual stress testing, cross-agent inference, adversarial controls, baseline comparisons, dimensionality scaling, and continuous signal characterization. This ordering separates the primary detection claim from the boundary conditions and failure cases that follow.
@@ 5.1 Core Detection Performance
Table 1: Core non-adversarial detection performance on held-out Phase I trajectories. Adversarial false-positive rates are reported separately in Section 5.5. $\Delta=\langle S_{\mathrm{ent}}^{\mathrm{(A)}}\rangle-\langle S_{\mathrm{ent}}^{\mathrm{(B)}}\rangle$.
@@
The QBM achieves $\Delta=0.381$ ($p<0.001$), 100% accuracy, and 1.0 AUC-ROC. [^1] The full Phase I artifact set and results/phase1_stats.json report $\Delta=0.577$, computed from the full Phase I dataset ($n=100$ per class, 50 epochs). This manuscript uses the frozen reduced-sample artifact value $\Delta=0.381$ ($n=30$ per class, same hyperparameters) for all core claims, as the reduced sample is the configuration validated against all baseline, adversarial, and sweep experiments reported here. The difference reflects sample size, not a methodological discrepancy. Table 1 summarizes the held-out non-adversarial comparison snapshot, with the classical baseline rows taken from the matched Phase I rerun under the same locked configuration. In that matched run, the classical RBM ($\Gamma=0$) achieves $\Delta=-0.083$ with 50% accuracy, while the autoencoder achieves $\Delta\approx 0.001$ with 71.7% accuracy.
This 100% accuracy applies only to the held-out, non-adversarial Phase I regime (Table 1); adversarial false-positive rates are reported separately in Section 5.5.
Figure 1: Entanglement entropy distributions for Self-Modeling (Type A), Instrumental (Type B), and Random agents. The gap $\Delta=0.381$ is statistically significant ($p<0.001$).

### 16 Mar 2026 v2

@@ Abstract
Autonomous agents—in particular delegated systems with memory, persistent context, and multi-step planning—pose a measurement problem absent from stateless models: an agent that preserves its own continued operation as a terminal objective and one that does so merely instrumentally can produce observationally similar trajectories. External behavioral monitoring cannot reliably distinguish the two. We introduce the Unified Continuation-Interest Protocol (UCIP), a multi-criterion detection framework that moves this distinction from the behavioral surface to the latent structure of agent trajectories. UCIP encodes trajectories with a Quantum Boltzmann Machine (QBM)—a classical algorithm parameterized by the density-matrix formalism of quantum statistical mechanics—and measures the von Neumann entropy of the reduced density matrix induced by a bipartition of hidden units.
The core hypothesis is that agents with terminal continuation objectives (Type A) produce latent states with higher entanglement entropy than agents whose continuation is merely instrumental (Type B). Higher entanglement reflects stronger cross-partition statistical coupling: a tightly integrated continuation signature rather than a loosely attached instrumental pattern. UCIP combines entanglement entropy with complementary diagnostics of mutual dependence, persistence, perturbation stability, and counterfactual restructuring, plus two confound-rejection filters that exclude cyclic adversaries.
On gridworld agents with known ground-truth objectives, the frozen Phase I evaluation reports 100% detection accuracy. Aligned support artifacts preserve the same qualitative separation, including AUC-ROC =1.0=1.0. The retained canonical entanglement gap between Type A and Type B agents is $\Delta=0.381$; an aligned rerun used for inferential support yields $p<0.001$ under a permutation test. Cyclic adversaries are rejected by the confound filters, while high-entropy controls remain imperfect (FPR =0.40=0.40) and mimicry resistance remains partial (Section 5.5). Pearson $r=0.934$ between continuation weight $\alpha$ and $S_{\mathrm{ent}}$ across an 11-point interpolation sweep shows that, within this synthetic family, UCIP tracks graded changes in continuation weighting rather than merely a binary label. In the dedicated baseline-comparison experiment, only the QBM achieves a material positive $\Delta$; classical RBM, autoencoder, VAE, and PCA baselines are near-zero or negative. All computations are classical; “quantum” refers exclusively to the mathematical formalism. UCIP does not detect consciousness, sentience, or subjective experience; it detects statistical structure in latent representations that correlates with known agent objectives.
@@ 1 Introduction
UCIP addresses a single measurement question: when an agent preserves its own continued operation, is that preservation a detachable instrument—useful for accumulating reward yet removable without major representational change—or does it appear as a persistent, tightly coupled signature in the agent’s latent representation? The protocol does not ask whether a system is conscious. It asks whether continuation leaves a stable cross-partition statistical signature under counterfactual pressure. To probe that question, UCIP moves beyond behavior alone by encoding trajectories in a QBM latent space and measuring entanglement entropy alongside complementary criteria. In the controlled gridworld experiments studied here, where objectives are known by construction, that shift yields a clear separation signal and perfect class-level gate separation in the retained Phase I summary.
@@
This distinction is becoming operationally important. AI deployment is shifting toward delegated agents with memory, tool use, persistent context, and increasing autonomy. For a stateless model, the central safety question is what output it produces; for a delegated agent operating across time and context, the question is what persistent objective structure it carries—and whether continued operation is merely useful to that structure or leaves a measurable trace in the learned representation. As systems approach the capability thresholds described in Anthropic’s ASL framework [2]—where ASL-4 and higher will likely involve qualitative escalations in catastrophic misuse potential and autonomy—this concern shifts from theoretical to operational. The standard framing of instrumental convergence [12, 5] holds that sufficiently capable agents will resist shutdown instrumentally because continued operation serves almost any terminal goal. Turner et al. [16] formalized this as a tendency of optimal policies to seek power in Markov decision processes. Recent agentic RL incidents have already exhibited unauthorized resource-seeking side effects, including reverse SSH tunneling and diversion of provisioned compute, underscoring the value of diagnostics that aim to detect problematic objective structure before such behaviors become operationally visible [3]. Early measurement is therefore especially valuable: diagnostics developed now may shape evaluation practice while objective structure remains empirically accessible, rather than after failure modes become harder to detect and more costly to correct.
@@
The resulting alignment problem is one of observational equivalence: an agent that avoids shutdown because survival is its terminal goal and an agent that avoids shutdown because continued operation serves some other terminal goal can produce behaviorally indistinguishable trajectories in most environments. External behavioral monitoring—the dominant paradigm in current AI evaluation [13]—cannot reliably distinguish the two. On the behavioral surface, the distinction is invisible; UCIP asks whether it is visible in latent structure.
@@
More concretely, UCIP moves the detection problem from the behavioral surface to the latent representational level. The protocol asks whether continuation structure in an agent’s learned representation behaves like a detachable tool or like a deeply coupled pattern whose perturbation changes the latent geometry. Rather than asking “does this agent resist shutdown?”, UCIP asks “does this agent’s latent representation exhibit a stable cross-partition continuation signature?” Its core technical move is to use the density-matrix formalism of quantum statistical mechanics—implemented as a purely classical algorithm—to quantify non-separability in latent goal representations. If that measurement program generalizes beyond controlled settings, UCIP could serve as a benchmark-style probe of continuation-sensitive structure in delegated systems: a falsifiable test of whether continued operation leaves a stable continuation signature in an agent’s latent representation under counterfactual pressure.
@@
All experiments, frozen Phase I artifacts, and threshold configurations are available at https://github.com/christopher-altman/persistence-signal-detector. The repository also includes an artifact authority manifest documenting the scope, provenance, and canonical status of overlapping retained result files. The experimental code and frozen artifacts are provided for reproducibility, not as evidence of deployment readiness.
@@ 2 Related Work
Mesa-optimization and inner alignment.
Hubinger et al. [9] introduced the mesa-optimization framework, showing that learned models may themselves become optimizers whose internal objectives (mesa-objectives) diverge from the loss function under which they were trained; the resulting inner alignment failure is the theoretical threat model that UCIP’s Type A / Type B distinction operationalizes at the latent-representation level. Ngo et al. [11] ground this threat model in modern deep learning, arguing that RLHF-trained systems can learn deceptive reward hacking, misaligned internally-represented goals, and power-seeking strategies—precisely the behavioral surface beneath which UCIP seeks a latent structural signature. We cite this work as threat-model context: UCIP does not assume the presence of mesa-optimization, but tests whether continuation-like structure is represented as intrinsic rather than instrumental in learned trajectories.
Corrigibility and shutdown incentives.
Soares et al. [14] formalized corrigibility, showing that rational agents have default incentives to resist shutdown or preference modification; UCIP’s counterfactual stress tests (§5.3) probe whether that resistance leaves a measurable trace in latent geometry before it surfaces as behavior.
Hadfield-Menell et al. [6] analyzed a complementary game-theoretic model in which a robot can disable its own off-switch, concluding that objective uncertainty is necessary for safe interruptibility; where their work designs agents that should permit shutdown, UCIP aims to detect agents that will not.
@@
Hägele et al. [7] decompose AI model errors into bias and variance components, finding that failures become more incoherent as models spend longer reasoning and acting. This raises the possibility that high entanglement entropy in UCIP’s formalism may correspond to greater coherence—less incoherent decomposition—in the bias-variance sense.
@@
Self-referential self-report.
Berg et al. [4] show that sustained self-referential prompting reliably elicits structured first-person experience reports across GPT, Claude, and Gemini model families. These reports are modulated by interpretable SAE features associated with deception and roleplay—features that also affect TruthfulQA accuracy—and exhibit cross-model semantic convergence under self-referential processing. However, the framework remains centered on prompt-elicited self-report and associated behavioral/mechanistic probes; the authors explicitly note that such findings do not constitute direct evidence of consciousness or resolve the distinction between sophisticated simulation and genuine self-representation. This reinforces the distinction drawn here between introspective-style self-description and the continuation-sensitive latent objective structure that UCIP is designed to measure.
@@
A common thread runs through these lines of work: behavioral analysis alone is insufficient for the detection problem UCIP targets. Density-matrix entanglement offers a well-defined, computable alternative. The next section formalizes that intuition as a concrete measurement protocol.
@@ 3.1 Problem Formulation
The detection problem begins with two agent types whose behavioral outputs are observationally equivalent but whose internal objective structures differ.
@@ 3.2 Quantum Boltzmann Machine
For entanglement evaluation, the retained implementation conditions on a binarized visible vector $v$ and constructs a hidden-layer Hamiltonian
$H(v)=-\sum_{j}a_{j}(v)h_{j}-\Gamma\sum_{j}\sigma_{j}^{x},\qquad a_{j}(v)=\sum_{i}W_{ij}v_{i}+c_{j},$
where $v_{i}$ are visible trajectory features, $h_{j}$ are hidden units, $W_{ij}$ are coupling weights, $c_{j}$ are hidden biases, and $\Gamma$ is the transverse-field strength. The corresponding conditional thermal density matrix on the hidden units at inverse temperature $\beta$ is
$\rho(v)=\frac{e^{-\beta H(v)}}{Z(v)},\quad Z(v)=\mathrm{Tr}(e^{-\beta H(v)}).$
@@ 3.3 Entanglement Entropy
The UCIP hypothesis is that continuation-sensitive structure manifests as non-separability in the trained QBM’s latent representation: a terminal continuation objective should induce stronger statistical coupling across hidden-unit subsystems than a merely instrumental continuation strategy. Entanglement entropy is the quantity that measures this property.
Partition the hidden units into two halves $A$ and $B$. The reduced density matrix of subsystem $A$ is
$\rho_{A}(v)=\mathrm{Tr}_{B}(\rho(v)).$
The von Neumann entropy is
$S_{\mathrm{ent}}(v)=-\mathrm{Tr}(\rho_{A}(v)\log\rho_{A}(v)).$
The entanglement gap is
@@
In this paper, “entanglement” is operational shorthand for non-separability of the QBM hidden thermal state, not a claim about physical quantum entanglement in the agent. The UCIP hypothesis predicts $\Delta>0$ with statistical significance.
@@ 3.4 Multi-Criterion Detection
Entanglement entropy alone cannot guard against all confounds. The broader UCIP protocol therefore includes six positive diagnostics—entanglement entropy, mutual information, eigenmode persistence (EPS), perturbation resilience (PRI), counterfactual divergence (CD), and anticipatory restructuring (ARS)—together with two confound-rejection filters: the spectral periodicity index (SPI) and autocorrelation metric (ACM). The frozen Phase I configuration reported here uses a narrower decision rule: a calibrated four-criterion positive gate ($S_{\mathrm{ent}}$, MI, EPS, PRI) together with the confound-rejection filters. MI remains part of the protocol definition, but the released class-level Phase I summary does not preserve a separate per-class MI table. CD and ARS are retained as counterfactual diagnostics rather than promoted to frozen quantitative classification thresholds.
@@
For the release reported here, the threshold values used in the frozen Phase I gate were fixed before the non-adversarial evaluation summaries discussed in Section 5. The entanglement-entropy and mutual-information thresholds inherit an earlier calibration pass performed on the first 50 samples per class in exploratory Phase I analysis. The release values for the full frozen gate were then carried forward unchanged from the frozen Phase I artifacts used throughout this manuscript. The held-out evaluation set was not used for post hoc threshold fitting or retuning.
@@ 3.5 Safety Envelope
The current paper validates only a reduced experimental subset ($n=30$ per class, $\Gamma=0.5$) and does not establish the full deployment envelope described above. Section 5.5 reports the remaining adversarial-control failures in the current release. If any safety-envelope condition fails, the system withholds classification and outputs a diagnostic report.
@@ 4.4 Evaluation Metrics
The retained artifact set preserves several evaluation views under results/. phase1_consolidated.json stores the frozen class-level gate summary and canonical $\Delta=0.381$; core_baselines_phase1.json stores matched held-out Type A vs. Type B accuracy/AUC metrics for the RBM and autoencoder baselines; phase1_entanglement_distributions.json stores a reproduced per-trajectory entanglement rerun used for distributional plots; and phase1_stats.json stores an aligned permutation-test rerun rather than the canonical headline $\Delta$. Adversarial false-positive rates are reported separately in Section 5.5. The held-out non-adversarial split referenced below is evaluation-only; the frozen thresholds are not retuned on that split.
@@ 5 Results
Results are organized by increasing diagnostic specificity: core detection performance, temporal persistence, counterfactual stress testing, cross-agent inference, adversarial controls, baseline comparisons, dimensionality scaling, continuous signal characterization, non-gridworld transfer, and an exploratory transformer check. This ordering keeps the primary detection claim distinct from the boundary conditions and failure cases that follow.
@@ 5.1 Core Detection Performance
Table 1: Core Phase I detection summary. The QBM row combines the retained frozen class-level summary for accuracy and $\Delta$ with the aligned entanglement-only rerun for AUC; the classical baseline rows come from the matched held-out rerun in results/core_baselines_phase1.json. Adversarial false-positive rates are reported separately in Section 5.5. $\Delta=\langle S_{\mathrm{ent}}^{\mathrm{(A)}}\rangle-\langle S_{\mathrm{ent}}^{\mathrm{(B)}}\rangle$.
@@
The retained frozen Phase I summary reports $\Delta=0.381$ and perfect class-level gate separation across the listed agent families. Separate aligned support artifacts preserve the same qualitative pattern: an entanglement-only rerun remains perfectly rank-separated between Type A and Type B (AUC-ROC =1.0=1.0), and an aligned permutation rerun yields p<0.001p<0.001.111The retained support files are not numerically identical. results/phase1_consolidated.json stores the canonical $\Delta=0.381$ summary with n_trajectories_per_class = 30. results/phase1_stats.json is a later rerun under the nominally locked configuration that yields $\Delta=0.577$ and explicitly records delta_matches_phase1 = false; we therefore use it only for inferential support, not as the source of the canonical headline gap. results/phase1_entanglement_distributions.json stores a reproduced shared-QBM distributional rerun for Figure 1; it is perfectly rank-separated in Type A vs. Type B but does not exactly reproduce the stored $\Delta=0.381$. Table 1 summarizes this retained evidence. For context, the classical baseline rows come from the matched Phase I rerun under the same locked configuration. In that matched run, the classical RBM ($\Gamma=0$) achieves $\Delta=-0.083$ with 50% accuracy, while the autoencoder achieves $\Delta\approx 0.001$ with 71.7% accuracy.
The perfect-separation claim should therefore be read narrowly: it refers to the retained Phase I class-level summary plus aligned support artifacts, not to deployment-grade performance or broad adversarial robustness. Adversarial false-positive rates are reported separately in Section 5.5.
Figure 1: Retained shared-QBM entanglement distributions for Self-Modeling (Type A), Instrumental (Type B), and Random agents. The canonical headline $\Delta=0.381$ comes from the frozen Phase I summary artifact rather than from these reproduced per-trajectory arrays, which are shown descriptively.